# MMSFormer: Multimodal Transformer for Material and Semantic Segmentation

### 17 Jan 2024 v3

@@ 1 Introduction
In this paper, we propose a novel fusion block that can fuse information from diverse combination of modalities. We also propose a new model for multimodal material and semantic segmentation tasks that we call MMSFormer. Our model uses transformer based encoders [47] to capture hierarchical features from different modalities, fuses the extracted features with our novel fusion block and utilizes MLP decoder to perform multimodal material and semantic segmentation. In particular, our proposed fusion block uses parallel convolutions to capture multi-scale features, channel attention to re-calibrate features along the channel dimension and linear layer to combine information across multiple modalities. Such a design provides a simple fusion block that can handle an arbitrary number of input modalities and combine information effectively from different modality combinations.
To evaluate our proposed MMSFormer and fusion block, we focus on multimodal material segmentation on MCubeS [29] dataset and multimodal semantic segmentation on FMB [33] and PST900 [40] datasets. MCubeS dataset consists of four different modalities: RGB, angle of linear polarization (AoLP), degree of linear polarization (DoLP) and near-infrared (NIR). FMB dataset includes RGB and infrared modalities, while PST900 dataset comprises RGB and thermal modalities. A series of experiments highlight the ability of the proposed fusion block to effectively combine features from different modality combinations, resulting in superior performance compared to current state-of-the-art methods. Ablation studies show that different input modalities assist in identifying different types of material classes. Furthermore, as we add new input modalities, overall performance increases gradually highlighting the ability of the fusion block to incorporate useful information from new modalities.
@@ 3.3 Multimodal Fusion Block
After the linear fusion layer, we added a module for capturing and mixing multi-scale features. The module consists of two linear projection layers having parallel convolution layers in between them. Specifically, we apply $3\times 3$, $5\times 5$, and $7\times 7$ convolutions to capture multi-scale features and add them with the residual feature. We apply two linear projection layers, one before and one after the parallel convolutions, to mix information in the feature map. Ablation studies show that $3\times 3$, $5\times 5$, and $7\times 7$ convolutions are sufficient for optimal model performance. These steps can be performed as
@@ 4 Experiments and Results
Table 1: Performance comparison on Multimodal Material Segmentation (MCubeS) dataset [29]. Here A, D, and N represent angle of linear polarization (AoLP), degree of linear polarization (DoLP), and near-infrared (NIR) respectively. Method Modalities % mIoU DRConv [3] RGB-A-D-N 34.63 DDF [58] RGB-A-D-N 36.16 TransFuser [38] RGB-A-D-N 37.66 DeepLabv3+ [4] RGB-A-D-N 38.13 MMTM [23] RGB-A-D-N 39.71 FuseNet [18] RGB-A-D-N 40.58 MCubeSNet [29] RGB-A-D-N 42.46 CMNeXt [50] RGB-A-D-N 51.54 MMSFormer (Ours) RGB-A-D-N 53.11
Table 2: Performance comparison on FBM [33] dataset. We show performance for different methods from already published works. Methods Modalities % mIoU GMNet [63] RGB-Infrared 49.2 LASNet [25] RGB-Infrared 42.5 EGFNet [11] RGB-Infrared 47.3 FEANet [8] RGB-Infrared 46.8 DIDFuse [57] RGB-Infrared 50.6 ReCoNet [21] RGB-Infrared 50.9 U2Fusion [48] RGB-Infrared 47.9 TarDAL [32] RGB-Infrared 48.1 SegMiF [33] RGB-Infrared 54.8 MMSFormer (Ours) RGB-Infrared 61.7
@@ Methods
4.1 Datasets
4.2 Implementation Details
We utilized a polynomial learning rate scheduler with a power of 0.9 to dynamically adjust the learning rate during training. The first 10 epochs were designated as warm-up epochs with a learning rate of 0.1 times the original rate. For loss computation, we used the cross-entropy loss function. Optimization was performed using the AdamW [35] optimizer with an epsilon value of $10^{-8}$ and weight decay set to 0.01.
4.3 Performance Comparison with Existing Methods
We conducted a rigorous performance evaluation of our model compared to established baseline models for three datasets. The comprehensive results are summarized in Tables 46.
Results on MCubeS Dataset. Table 4 shows the overall performance comparison between our model and existing baseline models for MCubeS dataset. Our model achieves a mean intersection-over-union (mIoU) of 53.11%, surpassing the current state-of-the-art model. It shows 1.57% improvement over CMNeXt [50] and 10.65% improvement over MCubeSNet [29] models. To further analyze the performance of our model, we conducted a per-class IoU analysis and presented in Table 3. Our model performs better in detecting most of the material classes compared to the current state-of-the-art models. Notably, our model demonstrates a substantial improvement in the detection of human (+8.0%), plastic (+6.3%), cobblestone (+4.9%) and asphalt (+3.7%) classes while maintaining competitive or better performance in other classes. This led to the overall better performance and sets new state-of-the-art for this dataset.
4.4 Performance Comparison for Incremental Modality Integration

### 07 Apr 2024 v4

@@ 1 Introduction
In this paper, we propose a novel fusion block that can fuse information from diverse combination of modalities. We also propose a new model for multimodal material and semantic segmentation tasks that we call MMSFormer. Our model uses transformer based encoders [48] to capture hierarchical features from different modalities, fuses the extracted features with our novel fusion block and utilizes MLP decoder to perform multimodal material and semantic segmentation. In particular, our proposed fusion block uses parallel convolutions to capture multi-scale features, channel attention to re-calibrate features along the channel dimension and linear layer to combine information across multiple modalities. Such a design provides a simple and computationally efficient fusion block that can handle an arbitrary number of input modalities and combine information effectively from different modality combinations. We compare our fusion block with some of the existing fusion methods in terms of number of parameters and GFLOPs in Table 9.
To evaluate our proposed MMSFormer and fusion block, we focus on multimodal material segmentation on MCubeS [29] dataset and multimodal semantic segmentation on FMB [33] and PST900 [40] datasets. MCubeS dataset consists of four different modalities: RGB, angle of linear polarization (AoLP), degree of linear polarization (DoLP) and near-infrared (NIR). FMB dataset includes RGB and infrared modalities, while PST900 dataset comprises RGB and thermal modalities. We show the overall and per-class performance comparison in Table 4-5 for these datasets. A series of experiments highlight the ability of the proposed fusion block to effectively combine features from different modality combinations, resulting in superior performance compared to current state-of-the-art methods. Ablation studies show that different input modalities assist in identifying different types of material classes as shown in Table 8. Furthermore, as we add new input modalities, overall performance increases graduallyhighlighting the ability of the fusion block to incorporate useful information from new modalities. We summarize the results in Table 4 and 6 for FMB and MCubeS datasets respectively.
@@ 3.3 Multimodal Fusion Block
After the linear fusion layer, we added a module for capturing and mixing multi-scale features. The module consists of two linear projection layers having parallel convolution layers in between them. First we apply a linear transformation on $\hat{F}^{i}$ along the channel dimension by passing it through the first linear projection layer. It refines and tunes the features from different channels. Then we apply $3\times 3$, $5\times 5$, and $7\times 7$ convolutions to effectively capture diverse features across different spatial contexts. By employing convolutions with different sizes, the fusion block can attend to local patterns as well as capture larger spatial structures, thereby enhancing its ability to extract meaningful features from the input data. Finally we apply another linear transformation along the channel dimension using the second linear project layer to consolidate the information captured by the parallel convolutions, promoting feature consistency and enhancing the discriminative power of the fused features. These steps can be performed as
@@
We found that using 3 parallel convolution layers with sizes $3\times 3$, $5\times 5$, and $7\times 7$ provide optimal performance. Increasing the convolution kernel size reduces performance which we show in Table 7. As larger kernels reduce performance, we did not add more than 3 parallel convolutions in our model.
@@ 4 Experiments and Results
Table 1: Performance comparison on Multimodal Material Segmentation (MCubeS) dataset [29]. Here A, D, and N represent angle of linear polarization (AoLP), degree of linear polarization (DoLP), and near-infrared (NIR) respectively. Method Modalities % mIoU DRConv [3] RGB-A-D-N 34.63 DDF [59] RGB-A-D-N 36.16 TransFuser [38] RGB-A-D-N 37.66 DeepLabv3+ [4] RGB-A-D-N 38.13 MMTM [23] RGB-A-D-N 39.71 FuseNet [18] RGB-A-D-N 40.58 MCubeSNet [29] RGB-A-D-N 42.46 CBAM [46] RGB-A-D-N 51.32 CMNeXt [51] RGB-A-D-N 51.54 MMSFormer (Ours) RGB-A-D-N 53.11
Table 2: Performance comparison on FBM [33] dataset. We show performance for different methods from already published works. Methods Modalities % mIoU CBAM [46] RGB-Infrared 50.1 GMNet [64] RGB-Infrared 49.2 LASNet [25] RGB-Infrared 42.5 EGFNet [11] RGB-Infrared 47.3 FEANet [8] RGB-Infrared 46.8 DIDFuse [58] RGB-Infrared 50.6 ReCoNet [21] RGB-Infrared 50.9 U2Fusion [49] RGB-Infrared 47.9 TarDAL [32] RGB-Infrared 48.1 SegMiF [33] RGB-Infrared 54.8 MMSFormer (Ours) RGB-Infrared 61.7
4.1 Datasets
4.2 Implementation Details
We utilized a polynomial learning rate scheduler with a power of 0.9 to dynamically adjust the learning rate during training. The first 10 epochs were designated as warm-up epochs with a learning rate of 0.1 times the original rate. For loss computation, we used the cross-entropy loss function. Optimization was performed using the AdamW [35] optimizer with an epsilon value of $10^{-8}$ and weight decay set to 0.01. For CBAM [46], we use the same encoder, decoder and hyperparameters used in our experiments and replace our fusion block with CBAM [^1] https://github.com/luuuyi/CBAM.PyTorch module. To be specific, after extracting the feature maps from each input modality using the modality specific encoders, we add them (sum them up) and pass the combined feature map to the CBAM module.
4.3 Performance Comparison with Existing Methods
We conducted a rigorous performance evaluation of our model compared to established baseline models for three datasets. The comprehensive results are summarized in Tables 46. We report results for CBAM from our experiments. Other results are taken from published literature.
Results on MCubeS Dataset. Table 4 shows the overall performance comparison between our model and existing baseline models for MCubeS dataset. Our model achieves a mean intersection-over-union (mIoU) of 53.11%, surpassing the current state-of-the-art model. It shows 1.57% improvement over CMNeXt [51], 1.79% improvement over CBAM [46] and 10.65% improvement over MCubeSNet [29] models. To further analyze the performance of our model, we conducted a per-class IoU analysis and presented in Table 3. Our model performs better in detecting most of the material classes compared to the current state-of-the-art models. Notably, our model demonstrates a substantial improvement in the detection of plastic (+3.7%), fabric (+3.1%), asphalt (+2.3%), and cobblestone (2.3%) classes while maintaining competitive or better performance in other classes. This led to the overall better performance and sets new state-of-the-art for this dataset.
@@ Methods
Table 9: Comparison of number of parameters in the fusion block and GFLOPs for MCubeS dataset having 4 input modalities with an input shape of $(3\times 512\times 512)$ for each modality. Our fusion block shows significantly lower complexity compared to existing methods.
Methods Fusion Block Parameters (M) GFLOPs
CMNeXt [51] 16.63 6.47
MCubeSNet [29] 7.41 12.10
HRFuser [1] 1.72 17.50
CMX [50] 16.59 6.41
DDF (Resnet-101) [59] 28.10 4.10
MMSFormer (Ours) 3.23 2.47